OpenStudy (anonymous):

Determine whether the statement is true or false ,and explain why or give an example that shows it is false 1) every nth-degree polynomial has(n-1) critical numbers. 2) The maximum value of y=3sinx+2cosx is 5 3) arcsin^2x+arccos^2x=1

OpenStudy (anonymous):

4) If y=ax+b then \[\frac{ \Delta y}{ \Delta x }= \frac{ dy }{ dx }\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

first, lets look at the easy problem : #4

ganeshie8 (ganeshie8):

Notice that \(y = ax + b\) is in `slope-intercept` form : \(\text{slope} = a = \dfrac{y_2-y_1}{x_2 - x_1} = \dfrac{\Delta y}{\Delta x}\)

ganeshie8 (ganeshie8):

by definition, derivative = slope so \(\dfrac{\Delta y}{ \Delta x} = \dfrac{dy}{dx}\) is true whenever the function is a line.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about the other ones?

ganeshie8 (ganeshie8):

3) \(\arcsin^2x+\arccos^2x=1\) just give a counter example when \(x = 0\)

ganeshie8 (ganeshie8):

\(\arcsin^2(0)+\arccos^2(0)= 0^2 + \left(\frac{\pi}{2}\right)^2 \ne 1 \) Q.E.D.

OpenStudy (anonymous):

in the second one i should find the maximum then put it value in the function to find y, right ?

ganeshie8 (ganeshie8):

you can do that, or simply use the below formula : \(a\cos x + b\sin x = \sqrt{a^2+b^2} \cos (x - \alpha ) \)

ganeshie8 (ganeshie8):

\(\implies\) the max value of \(y=3\sin x+2\cos x \) is \(\sqrt{3^2 + 2^2}= \sqrt{13}\) so the given statement is false

OpenStudy (anonymous):

what about the first one ?

ganeshie8 (ganeshie8):

1) every nth-degree polynomial has(n-1) critical numbers. strictly speaking its false; for ex : \(P(x) = x^4\) has only one critical number : x = 0

ganeshie8 (ganeshie8):

below is a correct statement however : 1) every nth-degree polynomial has `at most` (n-1) critical numbers.

OpenStudy (anonymous):

@ganeshie8 thanks alot , and sorry for taking ur time :)

ganeshie8 (ganeshie8):

np.. . u wlc :)